OpenStudy (anonymous):

A car dealer advertises "All cars 19% off sticker price!" A buyer pays $15,930.95 for a car. Estimate the sticker price.

OpenStudy (anonymous):

16,000 is there any multiple choice

OpenStudy (anonymous):

No multiple choice

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (bibby):

19% (0.19) off of some price = 1- 0.19 = 0.81*that price 0.81x = 15,930.95

OpenStudy (anonymous):

I got your signal, Commissioner. Cost = (1 - 0.19)(Sticker price)

OpenStudy (anonymous):

Just as @bibby has.

OpenStudy (bibby):

Follow my father @douglaswinslowcooper 's advice. He's shown me the right path for many other things in life.

OpenStudy (anonymous):

haha me 2

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

@bibby is quite generous.

OpenStudy (tkhunny):

We're ESTIMATING, people! 19% isn't far from 20% 15930.95 isn't far from 16000. 100% - 20% = 80% 16,000 is 80% of 20,000 That's one way to estimate it. Feel free to use your imagination and find other effecgtive and fast ways to do so.

OpenStudy (anonymous):

true

OpenStudy (anonymous):

i said 16,000

OpenStudy (tkhunny):

16000 is a very poor estimate for the sticker price.

OpenStudy (anonymous):

so the estimate of the sticker price should be $20,000 ?

OpenStudy (anonymous):

thnx for the drop on my thoughts #suckatestimates

OpenStudy (anonymous):

maybe

OpenStudy (tkhunny):

You didn't read the last part, did you? It's an estimate! The estimate should be whatever you estimate! I demonstrated one estimate. If you like that one, and understand it, feel free to go with it. Or, make up your own estimate.

OpenStudy (anonymous):

Good point: "estimate." (4/5) sticker = 16000 Sticker = 20000

OpenStudy (bibby):

16000 is an estimate on the discounted car. We're finding the estimate of the sticker (original) value. see: above

OpenStudy (tkhunny):

The word "good" is probably not appropriate. Any estimate with a reasonable basis, that is sufficient for your needs is a "good" estimate.

OpenStudy (anonymous):

well, im will try to find the answer!! as i will good bye daddy(doug(LOL))

OpenStudy (anonymous):

So what is the answer here 20,000.00 ?

OpenStudy (tkhunny):

"All cars 19% off sticker price!" A buyer pays $15,930.95 for a car. Estimate the sticker price" You can improve your estimate, with just a little more effort, by keeping track of your errors - just a little. 19% is almost 20% - We will be estimating the discount a little too high. Thus, the implied Sale Price will be a little low. 15930.95 is almost $16000. We are estimating the sale price a little too high. Well, that's interesting. One estimate is a little high (16000) and the other a little low (100 - 20 = 80). Thus, our estimate is expected to be not far from the actual value. If they were both high or both low, this would be a different story.

OpenStudy (tkhunny):

You must pay attention to ALL of what you are seeing. What estimate did you produce? Which estimate do you understand? Produce an estimate and go with it. Justify your process and go with it. You are supposed to be learning to ESTIMATE. You shoudl NOT be focused on what you think someone else may believe is RIGHT or CORRECT. What estimate have you imagined? Do it!

OpenStudy (anonymous):

I agree with the estimates of 20% and 16,000.00 but then not sure where to go from here. This whole conversation kind of got off track

OpenStudy (anonymous):

ahha

OpenStudy (tkhunny):

No, it's fine. The rest is arithmetic. 20% is a good number because it is a lovely fraction, 1/5, and this convenience may simplify the arithmetic in the estimate. If you discount a car 20% (or 1/5), this leaves 80% (or 4/5) that you actually pay. Do you believe this?

OpenStudy (anonymous):

Yes

OpenStudy (tkhunny):

Well, then 16000 is 4/5 of the Sticker Price. Personally, I picked 4/5 because I knew 16000 was divisible by 4. If 16000 is 4/5 of the Sticker Price, what is 5/5 of the Sticker Price? It is an arithmetic problem. \(\dfrac{16000}{4/5} = \dfrac{16000}{4}\cdot 5 = 4000\cdot 5 = 20000\)

OpenStudy (anonymous):

Thanks I see it! Thanks for your help

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

lol